# LabelBench: A Comprehensive Framework for Benchmarking Adaptive Label-Efficient Learning

### 12 Jan 2024 v3

@@ 1 Introduction
To showcase the power of our framework, we conduct experiments that benchmark multiple deep active learning algorithms in combination with semi-supervised learning and large pretrained models; compared to existing active learning literature, our experiments yield state-of-the-art label-efficiency. To highlight some of our results, we observe a more than four-fold reduction (75% savings) in annotation cost over random sampling on CIFAR-10 (Figure 1(a)), a dataset known to be particularly challenging for active learning  [^1] As reported in seminal papers and common benchmarks such as Ash et al. (2019) (Figure 16), Ash et al. (2021) (Figure 10), Beck et al. (2021) (Figure 1) and Lüth et al. (2023) (Figure 6-8), they see less than two-fold reductions in annotation cost.. This improvement is further demonstrated in our experiments on ImageNet (Figure 1(b, c)). Under any fixed annotation budget, our experiments suggest active learning algorithms can consistently boost test accuracy by more than 1.2% and pool accuracy (accuracy of predictions on the pool of unlabeled training data, as defined in Section 4.2) by more than 5%. Compared to the previous best results in this setting Emam et al. (2021), our results yield at least 10% higher test accuracy. Overall, LabelBench provides a light-weight benchmarking framework for researchers to test their algorithms on under more realistic and large-scale scenarios.
@@ 2.1 Semi-supervised Training
In traditional supervised learning the model is only trained on the set of labeled examples, while in Semi-SL the model training is also informed by the remaining unlabeled examples in the pool. Intuitively, Semi-SL leverages the assumption that examples lying “nearby” to one another should belong to the same class, and therefore during training the model is encouraged to produce the same model output for these examples (for an overview of Semi-SL we refer the interested reader to Zhu (2005); van Engelen and Hoos (2020); Ouali et al. (2020)). Broadly speaking, modern Semi-SL methods implement this principle using a combination of Consistency Regularization — where model outputs of neighboring examples are regularized to be similar — and/or Pseudo Labeling — where unlabeled examples that the model is confident on are assigned artificial labels to supplement supervised training Sohn et al. (2020); Berthelot et al. (2020). In our pipeline, we implement FlexMatch Zhang et al. (2021), Unsupervised Data Augmentation Xie et al. (2020b), and Pseudolabeling Lee (2013).
Semi-supervised Training of Large Pretrained Models. The application of Semi-SL to fine-tuning large pretrained models is a nascent area of research. Cai et al. (2022) pioneered the application of Semi-SL methods to large-scale vision transformers by using a multi-stage pipeline of pretraining followed by supervised fine-tuning and finally semi-supervised fine-tuning. Lagunas et al. (2023) apply this pipeline to a fine-grained classification e-commerce task and demonstrate improved performance compared to standard supervised training. Semi-SL training on transformer architectures has also been successfully applied to video action recognition Xing et al. (2023). USB Wang et al. (2022b) is a benchmark that includes Semi-SL evaluations on large pretrained models such as ViT; however, it does not incorporate AL into its pipeline, as we do here.
@@ 3.1 Selection via Proxy
Table 1: Estimated cost of neural network training for ImageNet experiments when collecting 600,000 labels with $20$ iterations (batches of 30,000 labels per iteration). Here we display the total cost of running $12$ trials with CLIP ViT-B32 and FlexMatch Semi-SL training (Zhang et al., 2021). All AWS dollars are based on on-demand rates of EC2 P3 instances.
@@ 4 Benchmarking Active Learning Algorithms
To demonstrate the utility of our framework, we conduct experiments comparing popular deep AL strategies in combination with large pretrained models and semi-supervised training. Our results presented in Section 4.4 show better label efficiencies than state-of-the-art deep AL literature. Moreover, we discuss the accuracy gap by using selection-via-proxy under different settings.
@@ 4.1 Experiment Setup
    • Model training. At the beginning of each iteration, the dataset is partially labeled. We use Semi-SL techniques to fine-tune the vision transformer or train the proxy model from scratch. In particular, we experiment with Semi-SL techniques that minimize a supervised training loss on labeled examples and an unsupervised loss on unlabeled examples that uses pseudolabeling and/or some form of consistency regularization. Most of our experiments use FlexMatch (Zhang et al., 2021), but we also experiment with simpler methods such as Unsupervised Data Augmentation (UDA) (Xie et al., 2020b) and Pseudolabeling (Lee, 2013) to assess the sensitivity of our pipeline to the choice of Semi-SL technique.
@@ 4.3 Datasets
We first test on CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009) and ImageNet (Deng et al., 2009), all of which are standard datasets used in previous AL and Semi-SL papers. To further evaluate LabelBench on more realistic datasets, we also test on iWildCam Beery et al. (2021) and fMoW Christie et al. (2018), parts of the WILDS benchmark Koh et al. (2021). To the best of our knowledge, only a handful of existing studies, such as Tamkin et al. (2022); Mussmann et al. (2022); Bartlett et al. (2022), have evaluated label-efficient algorithms on these datasets, albeit under different experimental setups. The WILDS benchmark was originally intended to represent distribution shifts faced in the wild (i.e., OOD test sets); here we limit our evaluation to in-domain (ID) test set performance as an initial exploratory step. Using these datasets provides several advantages: 1) Both of them are highly imbalanced. 2) Fine-tuning pretrained large-scale models on them is more challenging than on ImageNet (e.g., ID test accuracy on fMoW is 73.3% Wortsman et al. (2022) when fine-tuning ViT-L14 end-to-end). 3) Unlike ImageNet and CIFAR10, whose examples are gathered by querying search engines with human validation, iWildCam and fMoW gather labels directly from human annotators, which aligns more closely with our pool-based active learning setting.
@@ 4.4 Results and Discussion
Importance of the Choice of Semi-SL We evaluate the effect of using alternative Semi-SL techniques for end-to-end finetuning. We compare FlexMatch against two common Semi-SL baselines, UDA (Xie et al., 2020a) and Pseudolabeling (Lee, 2013), on CIFAR-10/100 in Figure 6 (methods described in more detail in Appendix C). FlexMatch, the most sophisticated Semi-SL algorithm we implement, achieves higher accuracy than UDA and Pseudolabeling across all AL sampling strategies. On CIFAR-10, the differences are most pronounced in the early AL rounds but diminish during the later rounds. On CIFAR-100 however, there remains a considerable difference even at a 20% labeling budget. Notably, label efficiency seems to be affected more by the choice of Semi-SL method than by the choice of AL method. Similar to the generalization accuracies in Figure 6, we report pool accuracies in Figure 21.
@@
(a) CIFAR-10, Pseudolabeling
(b) CIFAR-10, UDA
(c) CIFAR-10, Flexmatch
(d) CIFAR-100, Psuedolabeling
(e) CIFAR-100, UDA
(f) CIFAR-100, Flexmatch
Figure 6: Generalization accuracies of different semisupervised learning algorithms when combined with AL and CLIP ViT-B32, with end-to-end finetuning. The results in (a),(b), and (c) are on CIFAR-10 while (d),(e), and (f) are on CIFAR-100.
@@ 5 Call for Contribution and Future Work
Trainer. Our experiments demonstrate the potential label savings provided by combining active learning with FlexMatch, UDA, and Pseudolabeling. To expand upon these results, it would be valuable to develop a benchmark of additional Semi-SL methods, evaluated in combination with AL and large pretrained models. To do so, one could instantiate specific Semi-SL trainer classes that inherit our template Semi-SL trainer.
Active Learning Strategy. As exhibited by our results, combining AL with Semi-SL and large pretrained models results in highly accurate and label-efficient models that demonstrate clear gains over passive learning. Therefore, we call on the AL community to evaluate their active selection algorithms under our more comprehensive and up-to-date benchmark. Additionally, besides several widely-used AL strategies we have implemented here, there exists a large suite of active selection methods in the literature that could potentially be implemented and evaluated Ren et al. (2021).
@@ 6 Conclusion
This work is supported by AFOSR/AFRL grant FA9550-18-1-0166 and by the NSF under Grant Nos. IIS-2106937 and IIS-2148367.
@@ Appendix C Semi-Supervised Learning Strategies
  • Pseudolabeling (Lee, 2013): A pseudolabeling based semisupervised learning method that assigns pseudolabels to unlabeled samples on which model confidence exceeds a fixed threshold.
  • UDA (Xie et al., 2020b): A consistency-regularization based semisupervised learning method that ensures that the model predictions are consistent on both weakly and strongly augmented versions of highly confident unlabeled samples.
  • FlexMatch (Zhang et al., 2021): A semisupervised learning method that uses both consistency-regularization (similar to UDA) and pseudolabeling on unlabeled samples. Unlike UDA and Pseudolabeling, this approach also uses a dynamic confidence threshold, dependent on both time and class, to select which unlabeled samples to use in the unsupervised loss.
@@ F.3 Learning a Shallow Neural Network
F.4 End-to-end Fine-tuning with different Semi-SL Methods
Here we evaluate the effect of using alternative Semi-SL techniques on the pool accuracy for the end-to-end finetuning. We compare FlexMatch against two common Semi-SL baselines, UDA and Pseudolabeling, on CIFAR-10/100 in Figure 21.
(a) CIFAR-10, Pseudolabeling
(b) CIFAR-10, UDA
(c) CIFAR-10, Flexmatch
(d) CIFAR-100, Psuedolabeling
(e) CIFAR-100, UDA
(f) CIFAR-100, Flexmatch
Figure 21: Pool accuracies of different semisupervised learning algorithms when combined with AL and CLIP ViT-B32, with end-to-end finetuning. The results in (a),(b), and (c) are on CIFAR-10 while (d),(e), and (f) are on CIFAR-100.
@@ F.5 Evaluation on Small Budgets
The standard AL setup in recent works Beck et al. (2021); Coleman et al. (2019) uses significantly larger labeling budgets than the standard Semi-SL setup (Sohn et al., 2020; Zhang et al., 2021; Lee, 2013; Xie et al., 2020b). In Table  3 and 4, we experiment with AL methods in the small budget setting and demonstrate that AL still demonstrates considerable gains in accuracy compared to random sampling.
@@ F.7 Alternative Semi-SL Methods with Selection-via-Proxy
In Table 10, 11, 12, 13, we assess how important the choice of Semi-SL algorithm is when we use SVP. We consider CIFAR10 and CIFAR100 for this comparison and for the proxy model consider Linear Probe.
@@
Table 10: Test Accuracy: Selection-via-proxy results of CIFAR10 using CLIP ViT-B32 for different Semi-SL algorithms. The results are evaluated with 10,000 labels. Confidence intervals are standard errors based on four trials.

### 01 Mar 2024 v4

@@ 1 Introduction
To showcase the power of our framework, we conduct experiments that benchmark multiple deep active learning algorithms in combination with semi-supervised learning and large pretrained models; our experiments reveal especially strong label-efficiency gains from active learning, demonstrating a significant difference from conventional beliefs in existing literature. To highlight some of our results, we observe a more than four-fold reduction (75% savings) in annotation cost over random sampling on CIFAR-10 (Figure 1(a)), a dataset that was believed to be particularly challenging for active learning  [^1] As reported in seminal papers and common benchmarks such as Ash et al. (2019) (Figure 16), Ash et al. (2021) (Figure 10), Beck et al. (2021) (Figure 1) and Lüth et al. (2023) (Figure 6-8), they see less than two-fold reductions in annotation cost.. This improvement is further demonstrated in our experiments on ImageNet (Figure 1(b, c)). Under any fixed annotation budget, our experiments suggest active learning algorithms can consistently boost test accuracy by more than 1.2% and pool accuracy (accuracy of predictions on the pool of unlabeled training data, as defined in Section 4.2) by more than 5%. Compared to the previous best results in this setting (Emam et al., 2021), our results yield at least 10% higher test accuracy. Overall, LabelBench provides a light-weight benchmarking framework for researchers to test their algorithms on more realistic and large-scale scenarios.
@@ 2.1 Semi-supervised Training
In traditional supervised learning the model is only trained on the set of labeled examples, while in Semi-SL the model training is also informed by the remaining unlabeled examples in the pool. Intuitively, Semi-SL leverages the assumption that examples lying “nearby” to one another should belong to the same class, and therefore during training the model is encouraged to produce the same model output for these examples (for an overview of Semi-SL we refer the interested reader to Zhu (2005); van Engelen and Hoos (2020); Ouali et al. (2020)). Broadly speaking, modern Semi-SL methods implement this principle using a combination of Consistency Regularization — where model outputs of neighboring examples are regularized to be similar — and/or Pseudo Labeling — where unlabeled examples that the model is confident on are assigned artificial labels to supplement supervised training (Sohn et al., 2020; Berthelot et al., 2020). In our pipeline, we implement Pseudolabeling  (Lee, 2013), Unsupervised Data Augmentation (Xie et al., 2020b), FlexMatch (Zhang et al., 2021a), FreeMatch (Wang et al., 2022c) and SoftMatch (Chen et al., 2023).
Semi-supervised Training of Large Pretrained Models. The application of Semi-SL to fine-tuning large pretrained models is a nascent area of research. Cai et al. (2022) pioneered the application of Semi-SL methods to large-scale vision transformers by using a multi-stage pipeline of pretraining followed by supervised fine-tuning and finally semi-supervised fine-tuning. Lagunas et al. (2023) apply this pipeline to a fine-grained classification e-commerce task and demonstrate improved performance compared to standard supervised training. Semi-SL training on transformer architectures has also been successfully applied to video action recognition (Xing et al., 2023). USB (Wang et al., 2022b) is a benchmark that includes Semi-SL evaluations on large pretrained models such as ViT; however, it does not incorporate AL into its pipeline, as we do here.
@@ 3.1 Selection via Proxy
Table 1: Estimated cost of neural network training for ImageNet experiments when collecting 600,000 labels with $20$ iterations (batches of 30,000 labels per iteration). Here we display the total cost of running $12$ trials with CLIP ViT-B32 and FlexMatch Semi-SL training (Zhang et al., 2021a). All AWS dollars are based on on-demand rates of EC2 P3 instances.
@@ 4 Benchmarking Active Learning Algorithms
To demonstrate the utility of our framework, we conduct experiments comparing popular deep AL strategies in combination with large pretrained models and semi-supervised training. Our results presented in Section 4.4 show significantly better label efficiencies than existing deep AL literature. Moreover, we discuss the accuracy gap by using selection-via-proxy under different settings.
@@ 4.1 Experiment Setup
    • Model training. At the beginning of each iteration, the dataset is partially labeled. We use Semi-SL techniques to fine-tune the vision transformer or train the proxy model from scratch. In particular, we experiment with Semi-SL techniques that minimize a supervised training loss on labeled examples and an unsupervised loss on unlabeled examples that uses pseudolabeling and/or some form of consistency regularization. Most of our experiments use FlexMatch (Zhang et al., 2021a), but we also experiment with simpler methods such as Unsupervised Data Augmentation (UDA) (Xie et al., 2020b) and Pseudolabeling (Lee, 2013) to assess the sensitivity of our pipeline to the choice of Semi-SL technique.
@@ 4.3 Datasets
We first test on CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009) and ImageNet (Deng et al., 2009), all of which are standard datasets used in previous AL and Semi-SL papers. To further evaluate LabelBench on more realistic datasets, we also test on iWildCam (Beery et al., 2021) and fMoW (Christie et al., 2018), parts of the WILDS benchmark (Koh et al., 2021). To the best of our knowledge, only a handful of existing studies, such as Tamkin et al. (2022); Mussmann et al. (2022); Bartlett et al. (2022), have evaluated label-efficient algorithms on these datasets, albeit under different experimental setups. The WILDS benchmark was originally intended to represent distribution shifts faced in the wild (i.e., OOD test sets); here we limit our evaluation to in-domain (ID) test set performance as an initial exploratory step. Using these datasets provides several advantages: 1) Both of them are highly imbalanced. 2) Fine-tuning pretrained large-scale models on them is more challenging than on ImageNet (e.g., ID test accuracy on fMoW is 73.3% (Wortsman et al., 2022) when fine-tuning ViT-L14 end-to-end). 3) Unlike ImageNet and CIFAR10, whose examples are gathered by querying search engines with human validation, iWildCam and fMoW gather labels directly from human annotators, which aligns more closely with our pool-based active learning setting.
@@ 4.4 Results and Discussion
Importance of the Choice of Semi-SL We evaluate the effect of using alternative Semi-SL techniques for end-to-end finetuning on CIFAR-10. We compare FlexMatch against two common Semi-SL baselines, UDA (Xie et al., 2020a) and Pseudolabeling (Lee, 2013), and two recent Semi=SL approaches, SoftMatch (Chen et al., 2023) and FreeMatch (Wang et al., 2023), on CIFAR-10 in Figure 6. FlexMatch, FreeMatch, and SoftMatch achieve higher accuracy than UDA and Pseudolabeling across all AL sampling strategies. On CIFAR-10, the differences are most pronounced in the early AL rounds but diminish during the later rounds. Notably, label efficiency seems to be affected more by the choice of Semi-SL method than by the choice of AL method. Similar to the generalization accuracies in Figure 6, we report pool accuracies in Figure 21.
@@
(a) Pseudolabel
(b) UDA
(c) FlexMatch
(d) SoftMatch
(e) FreeMatch
Figure 6: Generalization Accuracy on CIFAR-10 with Alternate Semi-SL algorithms. Each result is averaged over three trials with standard error shown as the confidence interval.
@@ 5 Call for Contribution and Future Work
Trainer. Our experiments demonstrate the potential label savings provided by combining active learning with FreeMatch, SoftMatch, FlexMatch, UDA, and Pseudolabeling. To expand upon these results, it would be valuable to develop a benchmark of additional Semi-SL methods, evaluated in combination with AL and large pretrained models. To do so, one could instantiate specific Semi-SL trainer classes that inherit our template Semi-SL trainer.
Active Learning Strategy. As exhibited by our results, combining AL with Semi-SL and large pretrained models results in highly accurate and label-efficient models that demonstrate clear gains over passive learning. Therefore, we call on the AL community to evaluate their active selection algorithms under our more comprehensive and up-to-date benchmark. Additionally, besides several widely-used AL strategies we have implemented here, there exists a large suite of active selection methods in the literature that could potentially be implemented and evaluated (Ren et al., 2021).
@@
Weak Supervision. Utilizing weak annotation sources to annotate is another growing subfield of label-efficient learning (Zhang et al., 2021b). Incorporating weak supervision with large pretrained models, Semi-SL and AL poses many open questions. We believe incorporating weak supervision in a rigorous way is an important next step in improving LabelBench.
@@ 6 Conclusion
This work is supported by AFOSR/AFRL grant FA9550-18-1-0166 and by the NSF under Grant Nos. CNS-2112471, IIS-2106937 and IIS-2148367.
@@ Appendix C Semi-Supervised Learning Strategies
  • Pseudolabeling (Lee, 2013): A pseudolabeling based semi-supervised learning method that assigns pseudolabels to unlabeled samples on which model confidence exceeds a fixed threshold.
  • UDA (Xie et al., 2020b): A consistency-regularization based semi-supervised learning method that ensures that the model predictions are consistent on both weakly and strongly augmented versions of highly confident unlabeled samples.
  • FlexMatch (Zhang et al., 2021a): A semi-supervised learning method that uses both consistency-regularization (similar to UDA) and pseudolabeling on unlabeled samples. Unlike UDA and Pseudolabeling, this approach also uses a dynamic confidence threshold, dependent on both time and class, to select which unlabeled samples to use in the unsupervised loss.
  • FreeMatch (Wang et al., 2022c): Like FlexMatch, FreeMatch is a semi-supervised learning method that uses both consistency-regularization (similar to UDA) and pseudolabeling on unlabeled samples. In addition to an adaptive threshold for assigning pseudo labels for every class, FreeMatch also utilize class-specific adaptive thresholds to encourse class diversity.
  • SoftMatch (Chen et al., 2023): While SoftMatch uses a consistency-regularization regime as UDA, FlexMatch and FreeMatch above, it uses a soft threshold technique in generating pseudo labels. Specifically, the small amount of high confidence examples are weighted higher while the vast amount of lower confidence examples are also “pseudo-labeled”, but weighted less aggressively.
@@ F.3 Learning a Shallow Neural Network
F.4 Additional Results for End-to-end Fine-tuning with different Semi-SL Methods
Here we evaluate the effect of using alternative Semi-SL techniques on the pool accuracy for the end-to-end finetuning on CIFAR10 in Figure 21. Furthermore, we include results for CIFAR100 on with Pseudolabeling, UDA, and FlexMatch in Figure 22.
(a) Pseudolabel
(b) UDA
(c) FlexMatch
(d) SoftMatch
(e) FreeMatch
Figure 21: Pool Accuracy on CIFAR-10 with Alternate Semi-SL algorithms. Each result is averaged over three trials with standard error shown as the confidence interval.
(a) Test Acc, Pseudolabeling
(b) Test Acc, UDA
(c) Test Acc, Flexmatch
(d) Pool Acc, Psuedolabeling
(e) Pool Acc, UDA
(f) Pool Acc, Flexmatch
Figure 22: Results on CIFAR100 with different Semi-SL algorithms. Similar to the results of CIFAR10, we find that the choice of Semi-SL algorithm is very salient.
@@ F.5 Evaluation on Small Budgets
The standard AL setup in recent works (Beck et al., 2021; Coleman et al., 2019) uses significantly larger labeling budgets than the standard Semi-SL setup (Sohn et al., 2020; Zhang et al., 2021a; Lee, 2013; Xie et al., 2020b). In Table  3 and 4, we experiment with AL methods in the small budget setting and demonstrate that AL still demonstrates considerable gains in accuracy compared to random sampling.
@@ F.7 Alternative Semi-SL Methods with Selection-via-Proxy
In Table 10, 11, 12, 13, we assess how important the choice of Semi-SL algorithm is when we use SVP. We consider CIFAR10 and CIFAR100 for this comparison and for the proxy model consider Linear Probe. SoftMatch and FreeMatch are only evaluated on CIFAR10.
@@
SoftMatch FreeMatch
Fine-tune Linear Probe Fine-tune Linear Probe
Confidence 97.97 $\pm$ .13 97.85 $\pm$ .03 97.95 $\pm$ .11 97.84 $\pm$ .13
Entropy 97.99 $\pm$ .14 97.91 $\pm$ .13 97.87 $\pm$ .12 97.98 $\pm$ .01
Margin 97.97 $\pm$ .04 97.85 $\pm$ .01 98.03 $\pm$ .05 97.90 $\pm$ .14
Coreset 97.97 $\pm$ .02 97.64 $\pm$ .07 97.75 $\pm$ .14 97.70 $\pm$ .05
GALAXY 97.87 $\pm$ .09 97.77 $\pm$ .01 97.97 $\pm$ .12 97.85 $\pm$ .10
BADGE 97.94 $\pm$ .11 97.83 $\pm$ .04 97.88 $\pm$ .09 97.83 $\pm$ .06
BAIT 97.97 $\pm$ .09 97.77 $\pm$ .14 97.90 $\pm$ .04 97.85 $\pm$ .02
Random 97.57 $\pm$ .05 97.75 $\pm$ .03 97.58 $\pm$ .09 97.73 $\pm$ .01
Best 97.99 $\pm$ .14 97.91 $\pm$ .13 98.03 $\pm$ .05 97.98 $\pm$ .01
Table 10: Test Accuracy: Selection-via-proxy results of CIFAR10 using CLIP ViT-B32 for additional Semi-SL algorithms. The results are evaluated with 10,000 labels. Confidence intervals are standard errors based on four trials.
@@
SoftMatch FreeMatch
Fine-tune Linear Probe Fine-tune Linear Probe
Confidence 99.94 $\pm$ .00 99.61 $\pm$ .00 99.94 $\pm$ .00 99.64 $\pm$ .01
Entropy 99.94 $\pm$ .00 99.63 $\pm$ .02 99.92 $\pm$ .03 99.60 $\pm$ .01
Margin 99.94 $\pm$ .00 99.63 $\pm$ .00 99.94 $\pm$ .00 99.63 $\pm$ .01
Coreset 99.49 $\pm$ .00 98.65 $\pm$ .05 99.47 $\pm$ .03 98.64 $\pm$ .01
GALAXY 99.90 $\pm$ .00 99.59 $\pm$ .00 99.89 $\pm$ .01 99.60 $\pm$ .01
BADGE 99.92 $\pm$ .00 99.60 $\pm$ .00 99.93 $\pm$ .00 99.57 $\pm$ .00
BAIT 99.62 $\pm$ .03 99.28 $\pm$ .03 99.62 $\pm$ .04 99.29 $\pm$ .02
Random 98.22 $\pm$ .00 98.21 $\pm$ .02 98.21 $\pm$ .04 98.24 $\pm$ .00
Best 99.94 $\pm$ .00 99.63 $\pm$ .00 99.94 $\pm$ .00 99.63 $\pm$ .01